If a student scored 77 points on a test where the mean score was 82 and the standard deviation was 3. The student's z score was
Answer

Answer by Theo(13342)

z = (x-m)/s
x = data value
m = mean
s = standard deviation.
in your problem:
x = 77
m = 82
s = 3
z = (77-82)/3 = -5/3 = -1.6666667 which rounds to -1.67
to find the area to the left of that z-score, you would go to the z-score table and look for -1.6 in the left hand column of that table which tells you what row you answer will be in. you then look for .07 in the top row of that table which tells you what column your answer will be in.
the intersection of column and row is the area to the left of the z-score of -1.67.
answer you should get if you look in the z-score table is .0475
this means that a z-score of -1.67 is better than 4.75% of the population z-score.
that's not very good.
here's the z-score table that i used.
http://lilt.ilstu.edu/dasacke/eco148/ztable.htm
this assumes that the distribution of scores is normal.